### v1

# Fourier PINNs: From Strong Boundary Conditions to Adaptive Fourier Bases

Abstract
Interest is rising in Physics-Informed Neural Networks (PINNs) as a mesh-free alternative to traditional numerical solvers for partial differential equations (PDEs). However, PINNs often struggle to learn high-frequency and multi-scale target solutions. To tackle this problem, we first study a strong Boundary Condition (BC) version of PINNs for Dirichlet BCs and observe a consistent decline in relative error compared to the standard PINNs. We then perform a theoretical analysis based on the Fourier transform and convolution theorem. We find that strong BC PINNs can better learn the amplitudes of high-frequency components of the target solutions. However, constructing the architecture for strong BC PINNs is difficult for many BCs and domain geometries. Enlightened by our theoretical analysis, we propose Fourier PINNs — a simple, general, yet powerful method that augments PINNs with pre-specified, dense Fourier bases. Our proposed architecture likewise learns high-frequency components better but places no restrictions on the particular BCs or problem domains. We develop an adaptive learning and basis selection algorithm via alternating neural net basis optimization, Fourier and neural net basis coefficient estimation, and coefficient truncation. This scheme can flexibly identify the significant frequencies while weakening the nominal frequencies to better capture the target solution’s power spectrum. We show the advantage of our approach through a set of systematic experiments.
1 Introduction
Physics-informed neural networks (PINNs) (Raissi et al., 2019) are innovative, mesh-free approaches for solving partial differential equations (PDEs). They offer computationally efficient alternatives to traditional mesh-based numerical methods such as finite elements and finite volumes (Reddy, 2019). The optimization of PINNs involves softly constraining neural networks (NNs) through customized loss functions designed to adhere to the governing equations of physical processes. Researchers have successfully applied PINNs in various domains—for example, they have been used to simulate the radiative transport equation (Mishra & Molinaro, 2021), which is crucial for radio frequency chip and material design  (Chen et al., 2020; Liu & Wang, 2019). Further, cardiovascular flow modeling (Kissas et al., 2020) is another application, along with various fluid mechanics problems (Cai et al., 2021), and high-speed aerodynamic flow modeling (Mao et al., 2020), among many others (Raissi et al., 2020; Chen et al., 2020; Jin et al., 2021; Sirignano & Spiliopoulos, 2018; Zhu et al., 2019; Geneva & Zabaras, 2020; Sahli Costabal et al., 2020; Sun et al., 2020; Fang & Zhan, 2019).
Despite these successes, the training of PINNs remains challenging in some instances. Recent studies have analyzed common failure modes of PINNs, particularly when modeling problems that exhibit high-frequency, multi-scale, chaotic, or turbulent behaviors (Wang et al., 2022b; 2020b; 2020a; a), or when the governing PDEs are stiff (Krishnapriyan et al., 2021; Mojgani et al., 2022). Rahaman et al. (2019) attributes the slow convergence to the high-frequency components of the target solution, identifying it as a “spectrum bias” in standard NNs, while learning low-frequency information from data is straightforward. Wang et al. (2020b) confirmed that this bias is also present in the PINN setting. These challenges often arise because applying differential operators over the NN in the residual complicates the loss landscape (Krishnapriyan et al., 2021). From an optimization perspective, Wang et al. (2020a) highlighted that the imbalance in gradient magnitudes between the boundary loss and residual loss (with the latter often being much larger) causes the residual loss to dominate training, leading to a poor fit to the boundary conditions. Wang et al. (2022b) confirmed this conclusion through a neural tangent kernel (NTK) analysis of PINNs with wide networks, finding that the dominant eigenvalues of the residual kernel matrix often result in the training primarily fitting the residual loss.
One class of approaches designed to mitigate the training challenges in PINNs involves setting different weights for the boundary and residual loss terms. For example, Wight & Zhao (2020) suggested incorporating a large multiplier for the boundary loss term to prevent the residual loss from dominating the training process. In contrast, Wang et al. (2020a) proposed a dynamic weighting scheme based on the gradient statistics of the loss terms, and Wang et al. (2022b) developed an adaptive weighting approach based on the eigenvalues of the NTK. Liu & Wang (2021) employed a mini-max optimization to update the loss weights via stochastic ascent, and McClenny & Braga-Neto (2020) used a multiplicative soft attention mask to dynamically re-weight the loss term for each data point and collocation point. To alleviate the spectrum bias in NNs, Tancik et al. (2020) randomly sampled a set of high frequencies from a large-variance Gaussian distribution to construct random Fourier features as input to the NN. Additionally, Wang et al. (2021) used multiple Gaussian variances to sample frequencies for Fourier features, aiming to capture multi-scale solution information within the PINN framework. While effective, the performance of this method is sensitive to the number and scales of the Gaussian variances, which are user-specified hyperparameters that are often difficult to optimize.
Another strategy to improve these challenges is to modify the NN architecture to exactly satisfy the boundary conditions (BCs) (Lu et al., 2021; Lyu et al., 2020; Lagaris et al., 1998; 2000; McFall & Mahan, 2009; Berg & Nyström, 2018; Lagari et al., 2020; Mojgani et al., 2022). We refer to these approaches as “strong BC PINNs”. Despite their effectiveness, these approaches face several limitations. They are usually limited to tasks with relatively simple and well-defined physics, requiring significant craftsmanship and complex implementations even in straightforward problem settings. Consequently, these strategies are less flexible than the original PINN framework, which employs a soft constraint approach for boundary condition satisfaction. Additional complications arise for challenging physical systems governed by invariances or conservation laws, such as energy or momentum conservation, due to often poorly understood and imprecisely defined physical laws. Incorporating these laws effectively into the NN architecture makes extending strong BC PINNs to handle more complex tasks difficult. Further, designing a custom ansatz requires tailoring the NN architecture for each specific boundary condition and domain, which is often impractical for complicated domains. However, when properly designed, these techniques can achieve highly accurate solutions.
In our work, we delve deeper into the training challenges of PINNs for learning high-frequency and multi-scale solutions. We analyze the mechanisms behind the success of strong BC PINNs and explore ways to incorporate these successful strategies into a more general PINN architecture. Our specific contributions are as follows:
  • We first examine a strong BC PINN architecture for simple Dirichlet boundary conditions as proposed by Lu et al. (2021). This variant integrates a fixed polynomial boundary function into the NN architecture to exactly satisfy the boundary conditions. While it shows significant improvements over the standard PINN, especially for higher frequency problems, it struggles to predict solutions with frequencies above a certain threshold due to its static nature. To address this limitation, we propose a new strong BC PINN architecture featuring an adaptive parameter optimized during training. This parameter adjusts the boundary function’s sharpness to match the true solution, thereby improving accuracy for higher-frequency solutions compared to the static polynomial variant.
  • We conduct a Fourier analysis on both strong BC PINNs compared to the standard PINN. Through the Fourier series convolution theory, we discovered that multiplying the NN by the strong boundary function significantly enhances the learning speed and accuracy of the higher frequency coefficients in the target solution. In contrast, standard PINNs struggle to capture coefficients in the high-frequency domain accurately. This analysis complements and confirms the aforementioned NTK work.
  • Inspired by our Fourier analysis, we develop Fourier PINNs. This novel PINN architecture enhances frequency learning within the true solution, comparable to strong BC PINNs, regardless of specific boundary conditions, domain, or underlying physical properties. The Fourier PINN architecture integrates a standard NN with a linear combination of Fourier bases, with frequencies uniformly sampled from an extensive pre-set range. We implement an adaptive learning and basis selection algorithm that alternately optimizes the NN basis parameters and the coefficients of the NN and Fourier bases while pruning insignificant bases. This approach efficiently identifies significant frequencies, supplements those missed by the NN, and improves frequency amplitude estimation (i.e., the basis coefficient) while maintaining computational efficiency. Unlike previous methods, this approach only requires specifying a sufficiently large range and small spacing for the Fourier bases without concern for the actual number and scales of frequencies in the true solution.
  • We evaluate Fourier PINNs on several benchmark PDEs characterized by high-frequency and multi-frequency solutions. In all cases, Fourier PINNs consistently achieve low solution errors (e.g., $\sim 10^{-3}$ or $\sim 10^{-4}$). In contrast, standard PINNs invariably fail to achieve comparable accuracy. PINNs with random Fourier features (RFF-PINNs) often fail across various Gaussian variances and scales, indicating high sensitivity to these parameters. Additionally, we test spectral methods, PINNs with large boundary loss weights (Wight & Zhao, 2020), and PINNs with adaptive activation functions (Jagtap et al., 2020). Fourier PINNs consistently outperform all these methods.
The remainder of this paper is structured as follows: Section 2 provides the necessary background and notation, while in Section 4, we present our Fourier analysis of strong BC PINNs and explain the success of the strong boundary ansatz methodology. Our new Fourier PINN architecture and its training routines are described in Section 5. Section 6 details our numerical experiments and findings, including assessments of computational cost and accuracy compared to baseline methods. Finally, Section 7 discusses the results and outlines future research directions.
2 Background
In this section, we first describe the general optimization problems we are addressing with physics-informed neural networks (PINNs) following the formulation presented in Raissi et al. (2019). We then discuss the specifics of strong boundary condition enforcement.
2.1 General Overview of Physics-Informed Neural Networks
The PINN framework uses NNs to estimate solutions to partial differential equations (PDEs). Consider a PDE of the following general form,
${\mathcal{F}}[u]({\bf x})=f({\bf x}),\;\;{\bf x}\in\Omega,$ (1)
where ${\mathcal{F}}$ is a linear or non-linear differential operator and $u$ represents the unknown solution and $\Omega$ is the problem domain in $\mathbb{R}^{d}$. The general boundary conditions are then,
${\mathcal{B}}[u]({\bf x})=g({\bf x}),\;\;{\bf x}\in\partial\Omega,$ (2)
where $\partial\Omega$ is the boundary of the domain and $\mathcal{B}$ is a general boundary-condition operator.
To solve the PDE, the PINN uses a deep NN, $u_{N}({\bf x};{\bm{\theta}})$, to approximate the true solution $u({\bf x})$. For clarity in later sections, we follow the convention of Cyr et al. (2020) and define the output of the NN with width $W$ as a linear combination of a set of non-linear bases such that,
$u_{N}({\bf x};\textbf{c},{\bm{\theta}}^{H})=\sum_{j=1}^{W}c_{j}\bm{\psi}_{j}({\bf x};{\bm{\theta}}^{H}),$ (3)
where each $\bm{\psi}_{j}$ are non-linear activation functions (such as Tanh) acting on the hidden layer outputs. Each $c_{j}$ for $j=1,...,w$ and ${\bm{\theta}}^{H}$ are the weights and biases in the last layer of the NN and hidden layers, respectively. Therefore, ${\bm{\theta}}=\{\textbf{c},{\bm{\theta}}^{H}\}$ form the set of all network parameters. Then, finding the optimal network parameters ${\bm{\theta}}^{*}$ involves minimizing the following composite loss function between boundary and residual loss terms,
${\bm{\theta}}^{*}=\operatorname*{arg\,min}\nolimits_{\bm{\theta}}\;\;\lambda L_{b}({\bm{\theta}})+L_{r}({\bm{\theta}}).$ (4)
Here,
$L_{b}({\bm{\theta}})=\frac{1}{M}\sum_{i=1}^{M}\left(\mathcal{B}[u_{N}]({\bf x}_{b}^{i})-g({\bf x}_{b}^{i})\right)^{2},$ (5)
is the boundary loss to fit the boundary condition with Lagrange multiplier $\lambda$, and
$L_{r}({\bm{\theta}})=\frac{1}{N}\sum_{i=1}^{N}\left({\mathcal{F}}[u_{N}]({\bf x}_{r}^{i})-f({\bf x}_{r}^{i})\right)^{2},$ (6)
is the residual loss to fit the equation. We minimize equation 4 by sampling $N$ collocation points $\{{\bf x}_{r}^{i}\}_{i=1}^{N}$ from the domain $\Omega$ and $M$ points $\{{\bf x}_{b}^{i}\}_{i=1}^{M}$ from the boundary of the domain $\partial\Omega$, and iteratively modifying the network parameters through gradient descent.
2.2 Strong Boundary Condition PINNs
Equation 5 can approximately enforce various types of boundary conditions, including Dirichlet, Neumann, Robin, and periodic. However, prior analysis by Wang et al. (2020a; 2022b) suggests that the instability in training PINNs likely arises from the competition between the weakly enforced boundary loss (5) and the residual loss (6) terms during optimization. This competition likely occurs because both loss terms are minimized simultaneously during training. However, the optimization algorithm tends to prioritize the minimization of the residual loss, as it typically dominates the gradient of the combined loss function. Consequently, this can lead to a scenario where the residual loss converges effectively but at the expense of the boundary loss, which remains inadequately optimized. As a result, the boundary conditions may not be satisfied, leading to poor model performance on new, unseen data or data at the domain’s boundaries.
One solution to mitigate the competition between the loss terms is to design a surrogate model that inherently satisfies the boundary conditions. This approach typically works by modifying the network architecture to satisfy the boundary conditions exactly, thus eliminating the need for a boundary enforcement loss term and further reducing the computational cost. This method of exact imposition of boundary conditions has become a standard approach in the PINN literature and is especially prevalent for Dirichlet and periodic boundary conditions (Lu et al., 2021; Yu et al., 2022; Wang et al., 2024).
This work primarily focuses on Dirichlet boundary conditions to simplify the analysis and implementations while addressing a significant and common scenario in physical systems. Dirichlet boundary conditions provide a clear and straightforward framework to demonstrate the effectiveness of the surrogate model approach, as they require the solution to take specific values at the boundaries, which is relatively easy to enforce exactly on simplified domains. To illustrate, we consider the 1D Dirichlet boundary condition defined as,
$x\in[a,b],\;\;\;u(a)=u(b)=g(x).$ (7)
For this case, the surrogate model is typically defined as,
$u_{\bm{\theta}}(x)=g(x)+\phi(x)u_{N}(x),$ (8)
such that $u_{\bm{\theta}}(x)$ is the solution function, $g(x)$ is the provided boundary function, $u_{N}(x)$ is the output of the neural network, and $\phi(x)$ is a composite distance function that zeroes out on the boundaries, ensuring the boundary conditions are met without explicit penalty terms. By construction, $u_{\bm{\theta}}(a)=u_{\bm{\theta}}(b)=g(x)$, and thus the boundary condition in equation 7 is strictly and automatically satisfied. To estimate the parameters of $u_{\bm{\theta}}$, we only need to minimize the residual loss $L_{r}({\bm{\theta}})$. Most PINN surrogate model designs rely on distance functions based on the theory of R-functions (Sukumar & Srivastava, 2022), which are smooth mathematical functions that encode Boolean logic and facilitate the combination of simple shapes to form complex geometries (Rvachev & Sheiko, 1995). Some meshfree Galerkin methods have utilized the capability of R-functions to enforce boundary conditions smoothly and exactly to enhance the precision and robustness of numerical simulations for solving boundary-value problems (Shapiro & Tsukanov, 1999; Akin, 2014; Tsukanov & Posireddy, 2011).
Figure 1: This figure shows the polynomial distance function $\phi_{poly}(x)$ and the exponential distance function $\phi_{exp}(x)$ with different $\alpha$ values. The curves for $\phi_{exp}(x)$ correspond to different $\alpha$ parameters, demonstrating how the shape of $\phi_{exp}(x)$ “sharpens” around the domain boundaries as $\alpha$ increases.
In this work, we compare the standard PINN to two strong BC PINNs using different distance functions to enforce Dirichlet boundary conditions exactly. The first distance function we consider is the polynomial distance function proposed in (Lu et al., 2021), defined as,
$\phi_{poly}(x)=(x-a)(b-x).$ (9)
This function satisfies the boundary conditions by zeroing out at the endpoints $x=a$ and $x=b$. While Lu et al. (2021) demonstrates promising results using this formulation, the polynomial distance function $\phi_{poly}(x)$ remains static throughout the training process. The absence of adaptability limits the model’s performance across different problems and domains, whereas introducing adaptability would allow the model to adjust to the varying complexities within the solution space.
To address this limitation, we propose an adaptive distance function to introduce flexibility in enforcing the boundary conditions, which we define as:
$\phi_{exp}(x)=(1-e^{\alpha(a-x)})(1-e^{\alpha(x-b)}),$ (10)
where $\alpha$ is a parameter that can be pre-set or optimized during training. The exponential form of equation 10 allows the function to adjust dynamically to better conform to the unique characteristics of the problem domain, thus enhancing the enforcement of boundary conditions across different scenarios and potentially leading to more accurate and efficient training outcomes.
Figure 1 illustrates the behavior of the polynomial distance function $\phi_{poly}(x)$ to the exponential distance function $\phi_{exp}(x)$ with various values of $\alpha$. As $\alpha$ varies, $\phi_{exp}(x)$ adjusts its shape, showcasing its ability to conform to different problem domains and complexities. We expect this dynamic adaptability to enhance the model’s performance by more precisely enforcing boundary conditions across diverse scenarios. We affirm this through a numerical example in the next section.
3 Boundary Condition Pathologies in PINNs
Recent works have employed strong boundary enforcement strategies  (Lu et al., 2021; Yu et al., 2022) but offer few insights into the specific surrogate model design process. The design and implementation of the surrogate models are often ad-hoc, with no systematic approach to their architectural design. This complicates their application and limits their effectiveness across boundary conditions and physical problems, especially when dealing with complex physics and domains. Moreover, the standard PINNs frequently encounter difficulties with high-frequency and multi-scale solutions, leading to inaccurate predictions (Wang et al., 2020b). We show that while the strong BC PINN model improves solution quality for problems with higher frequency compared to the standard PINN, the solution quality nevertheless degrades with larger frequencies. In the following, we empirically demonstrate how the standard PINN solution quality degrades as the actual frequency of the solution increases in both a linear 1D Poisson problem and a non-linear 1D steady-state Allen-Cahn problem. Specifically, we consider the fabricated solution $u(x)=\sin(kx)$ for a range of frequencies $k\in[2,6,10,14,18,22,26,30,34]$ and $x\in[0,2\pi]$. Therefore, by varying $k$, we can examine the performance of the standard PINNs compared to the strong BC PINNs when the solution $u$ includes different frequency information.
Specifically, we examine the 1D Poisson equation given by,
$\Delta u$ $=f(x)$ (11)
$u(0)=u(1)$ $=0,$ (12)
and derive the forcing function $f(x)$ from the fabricated solution $u(x)=\sin(kx)$ which gives $f(x)=-k^{2}\sin(kx)$. We also consider the 1D Allen-Cahn equation defined as,
$u_{xx}+u(u^{2}-1)$ $=f(x)$ (13)
$u(0)=u(1)$ $=0,$ (14)
which gives the forcing function $f(x)=-k^{2}\sin(kx)+\sin(kx)(\sin^{2}(kx)-1)$. Both are subject to the Dirichlet condition in equation 7 where $a=0$ and $b=2\pi$. We test these problems using a standard PINN, which explicitly incorporates the weakly enforced boundary loss and residual loss terms. We compare these results to two strong BC PINNs employing different distance functions (8) and show that this method exhibits slower solution degradation with increased frequency. We used the same NN architecture for the strong BC PINN and the standard PINN, including two hidden layers, $100$ neurons per layer, and Tanh activation. Both methods used the same 10K collocation points randomly sampled from the domain. We report the averaged results for each method from five random trials.
Figure 2: Relative $\ell_{2}$ error in predicting the solution $u(x)=\sin(kx)$ as a function of frequency $k$. The plots compare the performance of the standard PINN, the strong BC PINN with polynomial boundary function $\phi_{poly}$, and the strong BC PINN with exponential boundary function $\phi_{exp}$ for (left) 1D Poisson and (right) 1D steady-state Allen-Cahn equations. The shaded regions represent the error variability.
Figure 2 shows that both of the strong BC PINNs given by equation 9 and equation 10 push the surrogate model to inhibit the high-frequency components resulting in higher accuracy solutions, however, once the frequency surpasses a certain threshold, the static polynomial distance function $\phi_{poly}$ fails to provide any additional benefit. Figure 2 shows that the strong BC PINN ($\phi_{exp}$) given by equation 10 obtains higher accuracy solutions than both the standard PINN and the strong BC PINN ($\phi_{poly}$). However, it similarly hits a threshold where the benefits of the strong boundary enforcement diminish, albeit at a much higher frequency than the strong BC PINN ($\phi_{poly}$).
4 Fourier Analysis of Strong BC PINNs
In this section, we explain why the strong BC PINNs outperform standard PINNs using Fourier analysis. Both NNs and PINNs often exhibit spectral bias, quickly capturing low-frequency information but struggling with high-frequency components (Basri et al., 2020; Rahaman et al., 2019; Xu et al., 2019; Wang et al., 2021). This bias means that while NNs can quickly grasp general patterns (low frequencies), they struggle with capturing finer details (high frequencies), leading to potential inaccuracies in the solution. The frequency spectrum of a PINN showing spectral bias would show large coefficients at low frequencies, indicating that the NN has effectively captured the low-frequency components of the target solution. In contrast, the coefficients for high-frequency components would decay slowly and exhibit heavy tails, indicating that the network is not effectively filtering out the high-frequency noise or inaccuracies. These tails suggest that the high-frequency components, even though not dominant, still have significant amplitudes. Overall, the frequency spectrum would highlight a pronounced low-frequency peak with a gradual and insufficient reduction in the magnitude of higher frequencies, reflecting the network’s difficulty in learning and accurately representing the high-frequency aspects of the solution.
Figure 3: The frequency spectrum of the learned solution surrogate (a) and the convolution operation in the strong BC PINN ($\phi_{poly}$) with $\omega=15$. In both cases, the ground-truth solution is $\sin(\omega x)$. The left graph shows how the standard and strong BC PINNs handle frequency components, while the right graph illustrates the impact of the convolution operation on reducing high-frequency noise.
Figure 3 shows the frequency spectrum from the discrete Fourier transform of a standard PINN trained to solve the 1D Poisson problem defined in equation 12 with the true solution $u(x)=\sin(\omega x)$ for $x\in[0,2\pi]$, where the solution frequency is $\omega=15$. Although the standard PINN identifies the target frequency seen by the large coefficient at $\omega=15$, it exhibits large coefficients for higher frequencies, leading to heavy tails and reduced accuracy. In contrast, the frequency spectrum of the strong BC PINN (Figure 3) demonstrates that it not only captures the target frequency as accurately as the standard PINN, but its higher frequency coefficients decay much faster. This rapid decay weakens or excludes the influence of unnecessary high frequencies, resulting in significantly improved accuracy. In simpler terms, while both networks can identify the main pattern in the data (the target frequency), the strong BC PINN is better at ignoring unnecessary noise (high-frequency components) which leads to a more accurate overall solution.
To analyze why the distance function ($\phi(x)$) after multiplied to the NN in equation 8 can help obtain better coefficients for high frequencies, we first represent the general boundary function $\phi(x)$ as an infinite Fourier series,
$\phi(x)=\phi_{\infty}(x)$ $=\sum_{n=-\infty}^{+\infty}\hat{\phi}[n]\cdot e^{inx}$ (15)
where $i$ indicates the imaginary part and $n$ are integers. We first analyze the polynomial distance function by substituting equation 9 for $\hat{\phi}$ in equation 15 giving,
$\hat{\phi}_{poly}[n]=\frac{1}{2\pi}\int_{2\pi}x(2\pi-x)\cdot e^{-inx}{\rm d}x\;=\;-\frac{2}{n^{2}}.$ (16)
Similarly, to compare against our proposed adaptive distance function, we substitute in equation 10 for $\hat{\phi}$ in equation 15 and obtain,
$\hat{\phi}_{exp}[n]=\frac{1}{2\pi}\int_{2\pi}(1-e^{-\alpha x})(1-e^{\alpha(x-2\pi)})\cdot e^{-inx}{\rm d}x\;=\;\frac{\alpha(1-e^{2\pi\alpha})e^{-2\pi\alpha}}{\pi(\alpha^{2}+n^{2})}.$ (17)
Here, we expand the boundary functions $\phi(x)$ to be periodic (i.e., duplicating its definition in $[0,2\pi]$ to other intervals) and the period is $2\pi$. We can now obtain the Fourier transforms of each $\phi_{\infty}(x)$ by evaluating,
$\hat{\phi}(\omega)=\sum\nolimits_{n=-\infty}^{+\infty}\hat{\phi}[n]\cdot\delta\left(\omega-{\frac{n}{2\pi}}\right)$ (18)
where $\delta(\cdot)$ is the Dirac delta function and $\omega$ is the continuous frequency. Specifically, each $\hat{\phi}(\omega)$ represents the Fourier transform of $\phi_{\infty}(x)$, meaning ${\mathcal{F}}^{-1}\left\{\hat{\phi}(\omega)\right\}=\phi_{\infty}(x)$ where ${\mathcal{F}}^{-1}$ denotes the inverse Fourier transform.
Since $|\hat{\phi}_{poly}[n]|\propto 1/n^{2}$, the amplitude of the frequencies of $\phi_{poly}(x)$ decay quadratically fast with the increase of the absolute frequency value. This rapid decay is beneficial for denoising as it effectively suppresses high-frequencies leading to smoother approximations. However, it may result in the loss of fine details, especially if important signal components are present at higher frequencies. Alternatively, $|\hat{\phi}_{exp}[n]|\propto\alpha/(\alpha^{2}+n^{2})$, indicating that the frequency amplitudes follow Lorentzian decay controlled by the decay rate $\alpha$. Lorentzian decay is less aggressive in reducing high-frequency components, which helps preserve fine details and sharp transitions in the signal. This flexibility is useful in tuning the decay based on signal characteristics, however, the slower decay rate may lead to less effective noise suppression.
We can now look at the frequency spectrum of the full surrogate model $u_{\bm{\theta}}(x)$ in equation 8. We use the convolution theorem (McGillem & Cooper, 1991) to obtain
$\hat{u}_{\bm{\theta}}(\omega)=\hat{\phi}*\widehat{u}_{N}=\int_{-\infty}^{\infty}\hat{\phi}(\omega-\nu)\;\widehat{u}_{N}(\nu)\;{\rm d}\nu$ (19)
where * denotes convolution, and $\widehat{u}_{N}$ is the Fourier transform of the NN. Suppose $\omega>0$, we know that $\hat{\phi}(\omega-\nu)$ is obtained by first reflecting $\hat{\phi}(\nu)$ about the y-axis, and then shifting the frequency $\omega$ left along the spectrum. Then, we integrate $\widehat{u}_{N}(\nu)$ weighted by $\hat{\phi}(\omega-\nu)$. The larger the frequency $\omega$, the more $\hat{\phi}$ is moved left to obtain $\hat{\phi}(\omega-\nu)$ resulting in a stronger weighting of the high-frequency components during integration. Since the frequency coefficients of $\hat{\phi}$ decay quadratically fast ($|\hat{\phi}[n]|\propto 1/n^{2}$), the larger the portion of the tail is used, and the smaller the integration result. Specifically, with $\phi$, the corresponding amplitude of $\hat{u}_{\bm{\theta}}(\omega)$ decrease fast when $\omega$ increases [^1] The same conclusion applies when we consider $\omega<0$.. See Figure 3 for an illustration. Thereby, during the course of training, the boundary function pushes the surrogate model in equation 8 to weaken irrelevant high-frequency components to reduce large tails and to better capture the frequency spectrum.
5 Fourier PINNs
Our Fourier analysis in the previous section revealed that the standard PINNs inefficiently capture the true amplitude of frequencies in the power spectrum of their predicted solutions. Specifically, the analysis showed that although standard PINNs could identify the true solution frequency, they failed to prune large and noisy frequencies, as indicated by the large amplitudes assigned at the tail end of the power spectrum in Figure 3. While the strong BC PINNs address this issue by multiplying the NN with a boundary function—which helps decay high-frequency amplitudes faster—their overall benefit remains limited. As shown in Figure 2, performance degrades as $k$ increases, indicating the boundary function’s inability to learn ultra-high frequency solutions. Additionally, the strong BC PINNs are restricted to specific boundary conditions, such as Dirichlet conditions in equation 7, limiting their applicability.
5.1 Fourier PINNs
To overcome these challenges, we introduce Fourier PINNs, a novel PINN architecture. Fourier PINNs seamlessly handle diverse boundary conditions, like the standard PINNs, while emphasizing the true solution frequencies during training—akin to the strong BC PINNs. Specifically, in order to flexibly yet comprehensively capture the frequency spectrum, we first introduce a set of dense frequency candidates, $\{\omega_{n}\}_{n=1}^{K}$, evenly sampled from the range $[1,K]$. From this set of frequency candidates, we define a set of trainable Fourier bases $u_{B}$ as,
$u_{B}({\bf x};\textbf{a},\textbf{b})=\sum_{n=1}^{K}a_{n}\cos(2\pi\omega_{n}{\bf x})+b_{n}\sin(2\pi\omega_{n}{\bf x}).$ (20)
The Fourier PINN architecture additively combines the outputs of both the NN from equation 3 and the Fourier bases from equation 20 to generate the augmented prediction $u_{F}$. Specifically, the full output of Fourier PINNs is,
$u_{F}({\bf x};{\bm{\theta}})=\sum_{j=1}^{W}c_{j}\bm{\psi}_{j}({\bf x};{\bm{\theta}}^{H})+\sum_{n=1}^{K}a_{n}\cos(2\pi\omega_{n}{\bf x})+b_{n}\sin(2\pi\omega_{n}{\bf x}).$ (21)
We define ${\bf w}=\{c_{1},...,c_{W},a_{1},...,a_{K},b_{1},...,b_{K}\}$ as the set of all basis coefficients, and ${\bm{\theta}}=\{{\bm{\theta}}^{H},{\bf w}\}$ form the set of all trainable network parameters.
While we present our method in the one dimensional input case, it is straightforward to extend to multiple dimensions. Consider the two dimensional input ${\bf x}=[x_{1},x_{2}]$ for an example. We use the same set of frequencies to construct Fourier bases for each input. Specifically, we construct
$\bm{\phi}(x_{1})=[\cos(2\pi\omega_{1}x_{1}),\;\sin(2\pi\omega_{1}x_{1}),\;\ldots,\;\cos(2\pi\omega_{K}x_{1}),\;\sin(2\pi\omega_{K}x_{1})],$
and
$\bm{\phi}(x_{2})=[\cos(2\pi\omega_{1}x_{2}),\;\sin(2\pi\omega_{1}x_{2}),\;\ldots,\;\cos(2\pi\omega_{K}x_{2}),\;\sin(2\pi\omega_{K}x_{2})].$
Then we apply the cross-product of $\bm{\phi}(x_{1})$ and $\bm{\phi}(x_{2})$ to obtain the tensor-product Fourier bases for ${\bf x}$. The surrogate model is given by,
$u_{\bm{\theta}}({\bf x})=u_{N}({\bf x})+\bm{\beta}^{\top}{\rm vec}\left(\bm{\phi}(x_{1})\bm{\phi}(x_{2})^{\top}\right)$ (22)
where ${\rm vec}(\cdot)$ denotes the vectorization, and $\bm{\beta}$ are the coefficients of the Fourier bases. Note that the tensor-product of bases grows exponentially with the problem dimension, quickly becoming computationally intractable. One potential solution is to generate more sparse Fourier bases, such as total-degree or hyperbolic cross bases. Another option is to re-organize the coefficients $\bm{\beta}$ into a tensor or matrix and introduce a low-rank decomposition structure to parameterize $\bm{\beta}$ to save computational cost (Novikov et al., 2015). These explorations are left for future work.
5.2 Adaptive Basis Selection Training Algorithm
Given that the ground-truth solution likely contains significantly fewer frequencies than the augmented candidate bases ($u_{B}$), and our previous analysis showed that models often struggle to prioritize learning necessary frequencies, we developed an adaptive learning and basis selection algorithm. This algorithm flexibly identifies meaningful frequencies while pruning the inconsequential ones—allowing the model to focus on learning and improving the bases that significantly contribute to the solution by inhibiting the learning of unnecessary (and likely noisy) frequencies.
To aid in identifying the significant frequencies, we add an $L_{2}$ regularization term to the basis parameters ${\bf w}$. This additional regularization penalizes large coefficients for the basis functions, thereby promoting sparsity and encouraging the model to focus on learning the most significant frequencies. The full loss function we use is similar to PINNs (see equation 4), but with the addition of the $L_{2}$ regularization term. Specifically, the loss function is,
$L({\bm{\theta}})$ $=L_{b}({\bm{\theta}})+L_{r}({\bm{\theta}})+\frac{\alpha}{2}\|{\bf w}\|^{2},$ (23)
where $\|{\bf w}\|^{2}$ represents the $L_{2}$ regularization of ${\bf w}$ and $\alpha$ is the regularization strength defined by the user prior to initiating training. This regularization term helps prune or inhibit unnecessary frequencies, allowing the model to focus more on learning and improving the bases that significantly contribute to the solution.
Our adaptive basis selection algorithm optimizes the regularized loss function in equation 23 by building upon the hybrid least squares gradient descent method proposed by Cyr et al. (2020). Cyr et al. (2020) formulates the output of a neural network as a linear combination of non-linear basis functions, as shown in equation 3. Their algorithm alternates between optimizing the hidden weights (${\bm{\theta}}^{H}$) via gradient descent and the final layer coefficients ($\mathbf{c}$) through the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\textbf{c}}\|A\textbf{c}-\textbf{b}\|_{\ell_{2}}^{2},$ (24)
such that $\textbf{b}_{i}={\mathcal{L}}[u]({\bf x}^{i})$ for the data points $\{{\bf x}^{i}\}_{i=1}^{M}$, $A_{ij}={\mathcal{L}}[\psi_{j}({\bf x}_{i};{\bm{\theta}}^{H})]$ for $j=1,...,W$, and ${\mathcal{L}}$ is some linear operator. equation 24 extends to problems with multiple operators, such as those defined by equation 1 and equation 2 when ${\mathcal{F}}$ and ${\mathcal{B}}$ are linear operators.
Representing the output of Fourier PINNs in equation 20 as a linear combination of both adaptive bases (i.e., the hidden layers of the NN) and an augmented Fourier series ensures our model’s compatibility with the general approach of the alternating least squares and gradient descent optimization. While Cyr et al. (2020) hold the coefficients c constant while optimizing the bases ${\bm{\theta}}^{H}$ through gradient descent, we alternatively choose to jointly optimize both the basis coefficients ${\bf w}$ and the adaptive bases ${\bm{\theta}}^{H}$ during the gradient descent step. During the least squares step, we solve a similar problem to equation 24 modified to suit the Fourier PINN architecture. Specifically, we solve the least squares problem of the form,
$\operatorname*{arg\,min}\nolimits_{\bf w}\|A^{\prime}{\bf w}-\textbf{b}\|_{\ell_{2}}^{2},$ (25)
such that $\textbf{b}_{i}={\mathcal{L}}[u_{F}]({\bf x}_{i})$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$, and
$A_{ij}^{\prime}=\begin{cases}{\mathcal{L}}[\psi_{j}({\bf x}_{i};{\bm{\theta}}^{H})]&\text{for }j=1,...,W,\\ {\mathcal{L}}[\cos(2\pi\omega_{j}{\bf x}_{i})]&\text{for }j=W+1,...,W+K,\\ {\mathcal{L}}[\sin(2\pi\omega_{j}{\bf x}_{i})]&\text{for }j=W+K+2,...,W+2K.\end{cases}$ (26)
We define equation 25 for one operator ${\mathcal{L}}$, but the method extends to multiple operators. Additionally, we extend our least-squares method to non-linear operators, unlike many works that focus only on linear cases.
Specifically, when the operator ${\mathcal{L}}$ is nonlinear, the coefficient updates of ${\bf w}$ require more careful consideration. We consider non-linear operators ${\mathcal{L}}$, that can be decomposed into linear and nonlinear parts such as,
${\mathcal{L}}[\cdot]={\mathcal{G}}[\cdot]+{\mathcal{S}}[\cdot]$ (27)
where ${\mathcal{G}}$ is a linear operator and ${\mathcal{S}}$ is the nonlinear part. For illustration, consider the Allen-Cahn problem define by equation 14. We have ${\mathcal{G}}[u]=u_{xx}$ and ${\mathcal{S}}[u]=u(u^{2}-1)$, and solve equation 25 such that $\textbf{b}_{i}={\mathcal{L}}[u_{F}]({\bf x}_{i})-\textbf{s}_{i}$ for the data points $\{{\bf x}_{i}\}_{i=1}^{M}$. Here, $\textbf{s}_{i}={\mathcal{S}}[u_{F}]({\bf x})=u_{F}({\bf x}_{i};{\bf w})(u_{F}({\bf x}_{i};{\bf w})^{2}-1)$ are the Fourier PINN’s evaluations with the current ${\bf w}$ on the set of points, and
$A_{ij}^{\prime}=\begin{cases}{\mathcal{G}}[\psi_{j}({\bf x}_{i};{\bm{\theta}}^{H})]&\text{for }j=1,...,W,\\ {\mathcal{G}}[\cos(2\pi\omega_{j}{\bf x}_{i})]&\text{for }j=W+1,...,W+K,\\ {\mathcal{G}}[\sin(2\pi\omega_{j}{\bf x}_{i})]&\text{for }j=W+K+2,...,W+2K.\end{cases}$ (28)
In this case, since b is determined from the current ${\bf w}$ but fixed as a constant in the least squares estimation, the update of ${\bf w}$ is in essense a type of fixed point iteration. When ${\mathcal{L}}$ does not include any linear operator (which might be rare), we can use any continues optimization algorithm, e.g., L-BFGS, to update ${\bf w}$.
We integrate a custom basis pruning routine into the optimization algorithm to eliminate insignificant bases and refine model training. Our modified loss function defined in equation 23 enhances this hybrid optimization algorithm with an $L_{2}$ regularization term to identify and prioritize significant frequencies. Algorithm 1 outlines our adaptive learning routine. The algorithm begins with a joint optimization of ${\bm{\theta}}^{H}$ and ${\bf w}$ to provide a warm start to the model parameters. Next, we fix ${\bm{\theta}}^{H}$ while alternating between solving for the coefficients ${\bf w}$ using the least squares method and truncating bases with coefficients below a predefined threshold. Specifically, bases corresponding to $|w_{j}|\leq 10^{-3}$ are removed along with $w_{j}$. After completing the alternating solve and truncate steps for a set number of iterations, the algorithm jointly optimizes all remaining parameters and bases using the Adam optimization algorithm. We repeat this routine for a fixed number of iterations, followed by L-BFGS optimization until the final convergence.
Input: $\bm{\theta}^{H}_{0}$ (initialized hidden parameters)
Output: Optimized parameters $\bm{\theta}^{H}$ and weights $\bm{w}$
$\bm{w}\leftarrow\text{LS}(\bm{\theta}_{0}^{H})$ ;
  // Solve LS problem for $\bm{w}$
$\bm{\theta}^{H},\bm{w}\leftarrow\text{ADAM}(\bm{\theta}_{0}^{H},\bm{w})$ ;
  // Initialize $\bm{\theta}^{H}$ and $\bm{w}$
for $i=1,...$ do
       for $k=1,...$ do
             $\bm{w}\leftarrow\text{LS}(\bm{\theta}^{H})$ ;
              // Solve LS problem for $\bm{w}$
             if $|w_{j}|<\delta$ for each $w_{j}$ then
                   Prune the corresponding basis;
                   Delete $w_{j}$ from $\bm{w}$;
             end if
       end for
      $\bm{\theta}^{H},\bm{w}\leftarrow\text{ADAM}(\bm{\theta}^{H},\bm{w})$ ;
        // Jointly optimize all parameters using ADAM
end for
Run L-BFGS until convergence;
Algorithm 1 Adaptive Basis Selection Hybrid Least Squares/Gradient Descent
6 Experiments
In this section, we present a series of tests comparing our methods against a range of PDE problems. These include the 1D and 2D Poisson equations, which are canonical linear elliptic PDE problems, and the 1D and 2D (steady-state) Allen-Cahn equations, which are non-linear elliptic PDEs demonstrating the effect of adding a non-linear reaction term. Additionally, we analyze the 1D (one-way) Wave equation to observe the impact of incorporating a first time derivative. We first describe the baseline methods we compare against Fourier PINNs, then list our experimental details. Finally, we provide a detailed discussion after presenting the results comparing each baseline on the example problems.
6.1 Baselines
We benchmarked against the following state-of-the-art PINN models for solving high-frequency and/or multi-scale solutions:
  • (1)
    (PINN) standard PINNs as formulated in (Raissi et al., 2017);
  • (2)
    (RFF-PINN) Random Fourier Feature PINNs (Wang et al., 2021) with dynamically re-weighted loss terms derived from the NTK eigenvalues as in Wang et al. (2022b),
  • (3)
    (W-PINN) Weighted PINNs that down-weight the residual loss term to reduce its dominance and to better fit the boundary loss Wang et al. (2022b);
  • (4)
    (A-PINN) Adaptive PINNs with parameterized activation functions to increase the NN capacity and to be less prone to gradient vanishing and exploding Jagtap et al. (2020); and
  • (5)
    (Spectral) standard Spectral Methods (Boyd, 2001), which approximate solutions using a linear combination of Fourier bases, and estimates the basis coefficients via least mean squares [^2] Throughout the experiments, we used the same set of Fourier bases for the spectral method and our method..
We do not test against the conceptually similar Fourier Neural Operators Li et al. (2020) as they are designed to solve the inverse problem, thus requiring a data loss term. Our method is designed to solve forward problems and requires no training data within the domain.
Table 1: This table lists the number and corresponding scales of Gaussian variances utilized in the RFF-PINN. Scales were selected randomly from a set range of $[1,200]$. Preference was given to the subset $[1,20,50,100]$ to ensure a balanced distribution across the available range.
Number Scales
1 $(20)$,   $(50)$,   $(84)$,   $(100)$
2 $(1,50)$,   $(3,20)$,   $(19,71)$,   $(39,69)$,   $(50,100)$
3 $(44,47,165)$,   $(1,20,194)$,   $(20,50,100)$,   $(1,50,189)$,   $(38,112,119)$
[…]
6.2 Hyperparameters and experimental details
All code is implemented in PyTorch [^3] Code will be made available upon acceptance. For all the NN based methods, we used two hidden layers, with $100$ neurons per layer and the Tanh activation function. We ran every method for five random trials, and report the five number summaries for each test in boxplot form. For PINN, W-PINN, A-PINN and RFF-PINN, we first ran $40$K iterations of the Adam optimizer (Kingma & Ba, 2014) with a learning rate of $10^{-3}$, and then L-BFGS optimizer (Liu & Nocedal, 1989) until convergence (with the tolerance set to $10^{-9}$). In Fourier PINNs, we specified the range of frequency candidates in the Fourier bases by setting a maximum frequency $K$ and including all equally spaced frequencies in the set $\{1,2,...,K\}$. In Algorithm 1, we set $\alpha=0.1$, the number of iterations $T$ to 1K, the inner iteration number $A$ to 5, the truncation threshold $\delta$ to $10^{-4}$, and the outer iteration number $E$ to 40. RFF-PINNs require specifying both the number and scales of the Gaussian variances used to construct the random features. To comprehensively evaluate the effect of both hyperparameters, we tested RFF-PINNs with 20 different settings. We used one, two, three and five scales, and set the variances to common values suggested in (Wang et al., 2021), combined with randomly sampled ones. The exact sets of chosen scales are listed in Table 1. For W-PINN, we varied the weight of the residual loss from $\{10^{-1},10^{-3},10^{-4}\}$, and for A-PINN, we introduced a learned parameter for the activation function in each layer and update these parameters jointly.
Figure 4: 1D Poisson with $u(x)=\sin(100x)$
Figure 5: 1D Poisson with $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$
Figure 6: 1D Poisson with $u(x)=\sin(6x)\cos(100x)$
6.3 1D Numerical Experiments
In this section, we demonstrate the performance of our methods compared to each baseline on 1D problems including the 1D Poisson, 1D steady-state Allen-Cahn, and 1D One-Way Wave equation. We first tested a set of 1D Poisson problems (see equation 12) with the following ground-truth solutions,
$u(x)$ $=\sin(100x),$ (29)
$u(x)$ $=\sin(x)+0.1\sin(20x)+0.05\cos(100x),$ (30)
$u(x)$ $=\sin(6x)\cos(100x).$ (31)
The boundary conditions are derived from the true solution. We visualize the relative $L_{2}$ errors for each baseline method on these problems in Figures 4, 5, and 6 respectively. For Fourier PINNs, we set $K=150,200,250$ and $300$ for problems equation 29 and equation 30, then simply set $K=200$ for the remaining tests. Then, we examined the 1D steady-state Allen-Cahn equation of the form 14 with the same true manufactured solutions as for the 1D Poisson problem and report the results in Figures 7, 8, and 9. The final 1D problem we tested is the 1D One-Way Wave equation of the form,
$u_{t}+10u_{x}=v(x,t),\;\;x\in[0,1],t\in[0,1]$ (32)
$u(x=0,t)=u(x=1,t)=0$
$u(x,t=0)=\sin(\pi x)+\sin(2\pi x)$
with a true solution of,
$u(x,t)=\sin(\pi x)\cos(10\pi t)+\sin(2\pi x)\cos(20\pi t).$ (33)
The results are reported in Figure 10.
Figure 7: 1D Allen-Cahn equation with true solution $u(x)=\sin(100x)$
Figure 8: 1D Allen-Cahn with true multi-scale solution $u(x)=\sin(x)+0.1\sin(20x)+0.05\cos(100x)$
Figure 9: 1D Allen-Cahn equation with true solution $u(x)=\sin(6x)\cos(100x)$
Figure 10: 1D One-Way Wave equation with true solution $u(x,t)=\sin(\pi x)\cos(10\pi t)+\sin(2\pi x)\cos(20\pi t)$
Figure 11: 2D Poisson with $u(x,y)=\sin(100x)\sin(100y)$
Figure 12: 2D Poisson with $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
Figure 13: 2D Allen-Cahn with $u(x,y)=(\sin(x)+0.1\sin(20x)+\cos(100x))\cdot(\sin(y)+0.1\sin(20y)+\cos(100y))$
6.4 2D Numerical Experiments
In this section, we demonstrate the performance of our methods compared to each baseline on 2D problems, including the 2D Poisson and 2D steady-state Allen-Cahn equations. We first tested on the 2D Poisson problem given by,
$\Delta u$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (34)
with the following ground-truth manfactured solutions,
$u(x,y)$ $=\sin(100x)\sin(100y),$ (35)
$u(x,y)$ $=\sin(6x)\cos(20x)+\sin(6y)\cos(20y).$ (36)
We visualize the relative $L_{2}$ errors for each baseline method on these examples in Figures 11 and 12 respectively. We next tested on the 2D steady-state Allen-Cahn problem which is given by.
$u_{xx}+u_{yy}+u(u^{2}-1)$ $=f(x,y),\;\;x,y\in[0,2\pi],$ (38)
with the ground-truth solution,
$u(x,y)$ $=\left(\sin(x)+0.1\sin(20x)+\cos(100x)\right)$
$\;\;\;\cdot\left(\sin(y)+0.1\sin(20y)+\cos(100y)\right),$ (39)
and we report the results in Figure 13. Note that in these tests, the RFF-PINN method failed and resulted in large relative errors for all the single-scale settings, therefore we did not report these results.
6.5 Discussion
We can see that in all examples for both the 1D and 2D problems, the standard PINN, W-PINN, and A-PINN failed to obtain reasonable solutions, with large relative $L_{2}$ errors for all cases (e.g.,$\sim 1.0$). This implies that these methods are unable to capture high-frequency signals. Additionally, RFF-PINN can achieve small relative errors with particular scale settings; however, for most settings (60-70% of the 20 total RFF-PINN settings listed in Table 1), RFF-PINN failed with large solution errors. These results show that the success of RFF-PINN is very sensitive to the choice of number and scales, and there is no apparent correspondence between scale settings and problem solutions. For example, for the single-scale true solution (see Figure 4), the successful cases of RFF-PINN include two, three or even five scale settings, e.g.,(1, 20, 194), yet it failed with all the single-scale settings. The spectral method also failed to obtain good solutions, implying that only estimating a linear combination of Fourier bases are not sufficient. As a comparison, Fourier PINNs achieves reasonable solution accuracy in all the cases, with relative $L_{2}$ errors between $10^{-3}$ and $10^{-4}$. Our method is also robust to the choice of the frequency range $K$. As long as $K$ is large enough, our method nearly always selected the target frequency, and returned similar solution accuracy. Note that, to focus on the evaluation of the key idea, Fourier PINNs did not employ any loss term re-weighting schemes to further improve the accuracy, such as the NTK re-weighting used in RFF-PINN or mini-max updates. However, it would be straightforward to integrate these methods into the proposed algorithmic framework.
Figure 14: 2D Poisson equation with true solution $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
Figure 15: Predicted solutions (left) and absolute errors (right) of each method on the 2D Poisson equation with true solution $u(x,y)=\sin(6x)\cos(20x)+\sin(6y)\cos(20y)$
We finally visualize the solution of each method in solving a 2D Poisson equation (see Figure 12) to visualize the point-wise errors for each method within the two dimensional domain. The ground-truth solution is given in Figure 14 and the solution obtained by each method is shown in Figure 15. Note that for W-PINN and RFF-PINN, we report the result with the smallest relative $L_{2}$ error. While the spectral method also obtains good results, the point wise error biases much on the regions that are close to the boundary and the relative $L_{2}$ error is two orders of magnitude larger than Fourier PINN’s. RFF-PINN (with the best number and scale choice) can roughy capture the solution structure, but the accuracy is far worse than Fourier PINNs and the spectral method. For example, in Figure 12, RFF-PINN nearly failed with every choice of the number and scale set. We can see that overall, Fourier PINNs best recovers the original solution.
7 Conclusion
We have presented Fourier PINNs, a novel PINN extension to capture high-frequency and multi-scale solution information. Our adaptive basis learning and selection algorithm can automatically identify target frequencies and truncate useless ones, without the need for tuning the number and scales. Currently, we assume the spacing (granularity) of the candidate frequencies is fine-grained enough, which can be overly optimistic. In the future, we plan to develop adaptive spacing approaches to capture this granularity.
Acknowledgments
We acknowledge the use of large language models (LLMs) for manuscript editing.